# IncomeSCM: From tabular data set to time-series simulator and causal estimation benchmark

### v1

@@ 1 Introduction
Estimating causal effects of interventions is fundamental to improving decision-making by data-driven means. As a result, an ever-growing body of research develops machine learning algorithms for this task (Hernán and Robins, 2006; Shalit et al., 2017; Wager and Athey, 2018; Yoon et al., 2018; Künzel et al., 2019; Kennedy, 2023). However, identification of causal effects—and, therefore, evaluating these algorithms—relies either on intervention (e.g., randomized experiments) or on unverifiable assumptions (e.g., exchangeability) (Pearl, 2009). As experimentation is costly or prohibited in many domains, simulated environments are uniquely suited as benchmarks for causal effects estimation from observational data. They can be constructed to satisfy conditions for identifiability and to answer interventional queries, providing ground truth labels for estimates. But this strategy begs the question: How can we build simulators that reveal which causal estimators work best in the real world?
@@
Primarily, two kinds of data sets are used for offline evaluation of causal effect estimates: experimental data and (semi-) synthetic data. Experimental data (e.g., the LaLonde (Jobs) data set (LaLonde, 1986)) supports unbiased estimation by removing confounding by construction (Imbens and Rubin, 2015). But learning from unconfounded data is rarely the goal of the methods we seek to evaluate. To this end, researchers can introduced selection bias by subsampling observations depending on the intervention (Hill, 2011). A drawback of this approach is that experimental data tends to be small, both in sample size and dimensionality, and subsampling decreases its size even further. In other semi-synthetic data, e.g., IHDP (Hill, 2011), the analyst simulates one or more of the variables in the system of interest, typically the outcome variable.
@@
Simulators can be data-driven, learned from observational as in Medkit-Learn (Chan et al., 2021), or hand-designed as in the 2016 ACIC causal inference challenge (Dorie et al., 2019). Hand-designed simulators are great for designing benchmarks to test estimators in a particular challenge, such as learning from non-linear outcomes (Hill, 2011; Dorie et al., 2019), heteroskedastic noise (Hahn et al., 2019), or post-treatment variables. Their main drawbacks are that i) they scale poorly with the size of the system, requiring significant work and domain expertise by the designer, and ii) they often model overly simple or extreme systems, not reflecting the intricacies of real-world data (Kinyanjui and Johansson, 2022; Hernán, 2019). In contrast, purely data-driven simulators of potential outcomes can fit and generate complex relations between context, treatment, and outcome variables which are difficult to design by hand (Chan et al., 2021). However, they often lack structural constraints on the causal relationships between variables. As a result, they are insufficient for reasoning about alternative identification strategies, and for changing the nature of confounding or interventions (Hernán et al., 2019). Typically, they mimic the observational distribution of the chosen data set but do not support creating variants of the simulator where other parts of the system have been changed. The ability to generate multiple variants of simulated data is critical to stress test estimators.
@@ 2 Related work
Simulated data has long been used to evaluate estimators of causal effects as a way to get around the unavailability of counterfactual outcomes. The extent and nature of simulation varies from replacing a single variable, typically the outcome, with a simulated counterpart. A notable example, the “IHDP” data set was derived from a randomized study of the Infant Health and Development Program by removing a biased subset of participants and introducing a synthetic response surface (Hill, 2011). It was built to support multiple settings, see e.g., (Dorie, 2016), but the variant used by Shalit et al. (2017) has stuck as the primary benchmark. Generally, randomized data can be turned observational by subsampling cohorts with treatment selection bias, as in Twins (Louizos et al., 2017) or by adding observational data to an experimental cohort as in Jobs (Shalit et al., 2017). Both strategies benefit from context variables remaining untouched, representative of real-world data. A downside is that the sample size is fixed to the number in the original study. This limitation can be overcome by designing every variable in the system, such as in the 2016 data challenges of the Atlantic Causal Inference Conference (ACIC) (Dorie et al., 2019). However, this often leads to very simplistic simulators.
@@ 3 Constructing the simulator
We create IncomeSCM—a simulator of a sequential decision-making process associated with personal income. This will serve as an example of a general strategy for creating realistic and challenging benchmark tasks for causal effect estimation from cross-sectional observational data.
@@
A benefit of this strategy is that it blends the strengths of data-driven and knowledge-guided design: 1) Imposing a causal structure defines interventional distributions of all variables, allowing benchmarks to ask flexible causal queries of the simulator (e.g., about mediation or path-specific effects). 2) Fitting the initial states of variables adds realistic complexity to structural equations. 3) A sequential model creates causal effects of future outcomes that are interesting functions of early interventions and context variables even when transition mechanisms are hand-designed.
@@
and aim to estimate CATE from observational data generated by the simulator.
We use this idea in IncomeSCM-1.0, a simulator of the yearly progression of individuals’ demographics, work, studies and income, as recorded in the US census.
@@ 3.1 The base data set: Adult
IncomeSCM-1.0 is based on the widely-used Adult data set (Becker and Kohavi, 1996) and its release on the UC Irvine Machine Learning Repository. [^1] https://archive.ics.uci.edu/dataset/2/adult The set contains 48 842 observations of 14 variables and a sample weight, extracted from the 1994 US Census bureau database. The original task associated with the Adult set is to predict whether an individual will earn more than \$50 000 in a year and the full variable set comprises age, workclass, education (categorical and numeric), marital-status, occupation, relationship, race, sex, capital-gains/losses (summarized as capital-net in the simulator) hours-per-week (of work), native-country and income. We ignore the sample-weight in the simulator and drop any rows containing missing values in any column, leaving 30 162 samples. Statistics for these features are presented in Table 1.
@@
As primary outcome of interest $Y$, we use the yearly income of subjects in US dollars. In the Adult data set, income is coded as a binary variable, representing whether the yearly income of a subject exceeds \$$\$50000$, denoted here by $R\in\{0,1\}$. For IncomeSCM, we construct a continuous income variable $Y$ by first fitting a small random forest regressor $h_{R}$ to predict the original binary outcome $R$ as a function of all covariates $X$ (excluding the added variable studies). The continuous predictions $\hat{R}_{i}=h_{R}(X_{i})\in[0,1]$ of the regressor were computed for each subject $i$ and re-scaled by a constant $c$ to achieve a desired average income across the full cohort.
@@
For a causal estimation task to be realistic, the primary intervention of interest must be possible to implement in practice, for example, in an experiment (Hernán, 2019; Hernán et al., 2022). The Adult data set contains a small number of variables that could be directly intervened upon and serve as the primary intervention. Base variables age, sex, native-country, and race can be removed from consideration, as can capital-gain, capital-loss as these are outside of the control of any feasible experiment. occupation is feasible to change on a yearly basis, but most jobs are unavailable to most people. relationship and marital-status are possible to change, through e.g., divorce, but most people would not be willing to change these in order to optimize their income, let alone participate in an experiment to do so. workclass, occupation and hours-per-week are candidates that could be intervened upon a yearly basis. education is coded in Adult by the highest attained degree (e.g., Bachelor’s degree) or the number of years of study. However, in the time-scale of the simulator (one observation per year, $t=1,...,T$), an intervention of, say $A_{t}\leftarrow\mathrm{Doctorate}$ is not feasible. Instead, to showcase the blend of data-driven and knowledge-guided design, we create an intervention on studies, a constructed variable, with values “Full-time studies”, “Day course”, “Evening course” and “No studies”, representing the study activity of a person during a year. Studies directly affect the education level, current income and the propensity to study in the next year, and indirectly the future income of a subject.
@@ 3.2 The causal graph
We construct a cDAG for the Adult data set (the initial-state cDAG for IncomeSCM-1.0) by assigning direct causes to each variable. age, sex, race and native-country were taken to be base variables without causal parents in the data set. education was deemed to be causal of workclass, which in turn causes occupation and marital.status. marital-status and relationship are closely linked, and the causal direction was chosen to be from the former to the latter. Finally, hours-per-week and capital-net were determined by the previous variables, and Income is the furthest downstream, directly caused also by Studies, the hand-designed variable. The full graph is illustrated in Figure 1 (middle-left). A very similar graph was used by Nabi and Shpitser (2018) except, in their case, marital-status was a direct cause of education, not the other way around. In our case, relationship is a direct cause of transitions in studies.
@@ 3.3 The samplers
IncomeSCM implements structural equations through samplers, functions of the values of parent variables, and an independent source of noise. The variables in the Adult data set are either numeric or categorical. For numeric variables $V\in\mathbb{R}$, we use (with few exceptions) an additive noise model,
@@
where $f_{V}:\mathrm{\textbf{pa}}(V)\rightarrow\mathbb{R}$ is a model of the conditional mean, $\mathbb{E}[V\mid\mathrm{\textbf{pa}}(V)]$, and $\epsilon_{V}$ is a source of homoskedastic noise (e.g., Gaussian). For structural equations where observations of $(V,\mathrm{\textbf{pa}}(V))$ are available, $f_{V}$ is implemented by fitting a regression (e.g., a Random forest regressor) to observed data, and the scale of $\epsilon_{V}$ is chosen proportional to the regression error. capital-net is an exception to this form, since the majority of capital gains/losses is 0. For this, we blend a classifier which predicts non-zero values and a conditional regressor for the case where the value is non-zero.
@@
For equations with no available data, such as for studies, $f_{V}$ is designed by hand. In the case of studies, the initial-state distribution is modeled as a logistic regression with features and coefficients depending on the age of the subject, whether they are single parents, their current education level, and constant rates for different types of studies. For income1 and the transitions $g_{V}$ of most variables $V\in\mathcal{V}$, we use a blend of hand-designed rules and fitted regressions. For example, the sampler for relationshipt adopts a probability $p_{\mathrm{stay}}$, that the person will stay in their relationship, relationshipt = relationshipt-1. With probability $1-p_{\mathrm{stay}}$, relationshipt is resampled from the same data-driven model as the initial state, see Figure 1 (right).
@@ 4 Implementation & data set
Here, parents specify the direct causes of the initial state of the variable. sampler specifies the type and parameters of the sampler, in this case, the initial state of education is distributed according to a logistic regression of age, race, sex and native-country. seq_parents_curr specifies the direct causes of the transition probability from the current time step, and seq_parents_curr causes from the previous time step. seq_sampler specifies which function determines this probability. LogisticSampler is a general sampler, that outputs the probability of the given variable as a logistic-linear function of the direct causes specified in parents. General samplers can be reused for other data sets and are defined in samplers.py.
@@ 4.1 Benchmark data set
We construct a single-time causal effect estimation benchmark around the question: {mdframed}[backgroundcolor=verylightgray,linewidth=0.5pt,leftmargin=0pt,rightmargin=0pt,innertopmargin=1em,innerbottommargin=0.5em,roundcorner=3pt, linecolor=bordercolor] What is the effect on the income $Y$ of a subject at $t=5$, four years after a year of $A_{2}=$“Full-time studies” compared to $A_{2}=$‘No studies” at $t=2$, and then proceeding as they wish?
For this task, we create a cross-sectional observational data set by first sampling from the simulator at time $t=1,...,5$ where studies are selected by a “default” observational stochastic policy as defined by the samplers in IncomeSCM. Then, time $t=1$ is used to record the previous income level income_prev and studies studies_prev, prior to intervention. All other covariates are read off at time $t_{0}=2$. The binary intervention of interest is then performed at $t_{0}=2$, with $A_{2}=$ Full-time studies or $A_{2}=$ No studies. The outcome variable $Y_{T}=$incomeT is recorded at $T=5$. We use this to sample $n=50000$ samples with random seed $s=0$. First-order covariate statistics of covariates, interventions and outcomes for this seed can be seen in Table 1.
To create ground truth for causal effect estimates, we generate two analogous sets of trajectories where the same set of subjects $i=1,...,m$ are either all assigned $A^{i}_{2}=$Full-time studies or all assigned $A^{i}_{2}=$ No studies, thereby recording samples of both counterfactual outcomes $Y^{i}(0),Y^{i}(1)$ of these subjects. We control that the sets of subjects are identical before both interventions at $t_{0}=2$ by setting a shared seed, $s=1$. We use a different seed from the observational data set to measure out-of-sample generalization. The resulting data set(s) of observational and two interventional samples is referred to as IncomeSCM-1.0.CATE.
@@
    CATE, conditioned only on hours-per-week (not itself a valid adjustment set), binned uniformly every 10 years, using the minimal adjustment set from Task 2.
Valid adjustment sets can be derived by standard graphical criteria like the backdoor criterion (Pearl, 2009) from the edge list in Table 5. Note, as described above, that the observed intervention is performed at time $t=2$ with $t=1$ being the initial state. Thus, a minimal adjustment set for estimating the causal effect of intervening on studies2 is the set of direct causes of this variable. In IncomeSCM-1.0, studiest at times $t>0$ are directly caused by studiest-1, incomet-1, aget, sext,educationt, and relationshipt. [^2] In the implementation, the helper variable time is also passed to enable interventions at specific time steps. Expanding this set with any pre-intervention variables (such as other variables at times $t=1,2$) also results in valid adjustment sets. However, to estimate the CATE conditioned only on hours-per-week2, it is not sufficient to adjust only for this variable.
Table 1: Baseline characteristics in the IncomeSCM-1.0 cohort at $t=0$ and the base data set Adult (after removing missing values), as well as in-sample fit quality. Categorical variables are reported as number (rate) and continuous variables as mean (interquartile range). The proportion of subjects with income higher than \$50 000 is higher in IncomeSCM-1.0 than in Adult since the average income was calibrated to match the US population of 2023. Base variables do not involve modeling, only computing proportions/means. The capital-net regressor is fit only to sample with non-zero result. Fit quality is reported as $R^{2}$ for continuous variables and AUC for categorical variables.
@@
      United-States 9090 (90.9) 27504 (91.2)
sex = Female $3$ (33.3) 9782 (32.4) $\dagger$
@@
      White 8581 (85.8) 25933 (86.0)
      Black 951 (9.5) 2817 (9.3)
age 39.8 (31.0, 48.0) 38.4 (28.0, 47.0) $\dagger$
@@
      HS-grad 3271 (32.7) 9840 (32.6)
      Some-college 2139 (21.4) 6678 (22.1)
@@
      Prof-specialty 1349 (13.5) 4038 (13.4)
      Exec-managerial 1349 (13.5) 3992 (13.2)
@@
      Husband 3465 (34.6) 12463 (41.3)
      Wife 1343 (13.4) 1406 (4.7)
capital-net 1082.9 (0.0, 0.0) 1003.6 (0.0, 0.0)   0.11*
hours-per-week 42.1 (24.0, 59.0) 40.9 (40.0, 45.0) 0.28
@@
      Full-time studies 1524 (15.2)
      No studies 4989 (49.9)
income>50K (baseline) 4826 (48.3) 7508 (24.9)
income (baseline, \$1000) 64.5 (17.9, 90.4)
income (5 years, \$1000) 13.0 (99.8, 155.2)

### v2

@@ 1 Introduction
Estimating causal effects of interventions is fundamental to improving decision-making by data-driven means. As a result, an ever-growing body of research develops machine learning algorithms for this task (Hernán and Robins, 2006; Shalit et al., 2017; Wager and Athey, 2018; Yoon et al., 2018; Künzel et al., 2019; Kennedy, 2023). However, identification of causal effects—and, therefore, evaluation of these algorithms—rely either on interventions (e.g., randomized experiments) or unverifiable assumptions (e.g., exchangeability) (Pearl, 2009). As experimentation is costly or prohibited in many domains, simulated environments are uniquely suited as benchmarks for causal effects estimation from observational data. They can be constructed to satisfy conditions for identifiability and to answer interventional queries, providing ground truth labels for estimates. But this strategy begs the question: How can we build simulators that reveal which causal estimators work best in the real world?
@@
Primarily, two kinds of data sets are used for offline evaluation of causal effect estimates: experimental data and (semi-) synthetic data. Experimental data (e.g., the LaLonde (Jobs) data set (LaLonde, 1986)) supports unbiased estimation by removing confounding by construction (Imbens and Rubin, 2015). But learning from unconfounded data is rarely the goal of the methods we seek to evaluate. To overcome this, researchers can introduce selection bias synthetically by subsampling observations depending on the intervention, as in Twins (Louizos et al., 2017). A drawback of this approach is that experimental data tends to be small, both in sample size and dimensionality, and subsampling decreases its size even further. In other semi-synthetic data, e.g., IHDP (Hill, 2011), the analyst simulates one or more of the variables in the system of interest, typically the outcome variable.
@@
Simulators can be data-driven, learned from observational as in Medkit-Learn (Chan et al., 2021), or hand-designed as in the 2016 ACIC causal inference challenge (Dorie et al., 2019). Hand-designed simulators are great for designing benchmarks to test estimators in a particular challenge, such as learning from non-linear outcomes (Hill, 2011; Dorie et al., 2019) or heteroskedastic noise (Hahn et al., 2019). Their main drawbacks are that they i) scale poorly with the size of the system, requiring significant domain expertise by the designer, or that they ii) model overly simple or extreme systems, not reflecting the intricacies of real-world data (Kinyanjui and Johansson, 2022; Hernán, 2019). In contrast, purely data-driven simulators of potential outcomes can fit and generate complex relations between context, treatment, and outcome variables which are difficult to design by hand (Chan et al., 2021). However, they often lack structural constraints on the causal relationships between variables. As a result, they are insufficient for reasoning about different identification strategies, e.g., with instrumental variables, or with different adjustment sets, and for changing the nature of confounding or interventions (Hernán et al., 2019). In other words, mimicing only the observational distribution of a chosen system is not sufficient to stress test causal effect estimators.
@@ 2 Related work
Simulated data has long been used to evaluate estimators of causal effects as a way to get around the unavailability of counterfactual outcomes. The extent and nature of simulation vary from replacing a single variable (typically the outcome) with a simulated counterpart, to simulating an entire system of random variables. A notable example, the “IHDP” data set was derived from a randomized study of the Infant Health and Development Program by removing a biased subset of treated participants and introducing a synthetic response surface (Hill, 2011). It was built to support multiple settings, see e.g., (Dorie, 2016), but the variant used by Shalit et al. (2017) has stuck as the primary benchmark. Generally, randomized data can be turned observational by subsampling cohorts with treatment selection bias, as in Twins (Louizos et al., 2017) or by adding observational data to an experimental cohort as in Jobs (Shalit et al., 2017). Both strategies benefit from context variables remaining untouched, representative of real-world data. A downside is that the sample size is fixed to the number in the original study. This limitation can be overcome by designing every variable in the system, such as in the 2016 data challenges of the Atlantic Causal Inference Conference (ACIC) (Dorie et al., 2019). However, this often leads to very simplistic simulators.
@@ 3 Constructing the simulator
We create IncomeSCM-1.0—a simulator of a sequential decision-making process associated with personal income. This will serve as an example of a general strategy for creating realistic and challenging benchmark tasks for causal effect estimation from cross-sectional observational data.
@@
A benefit of this strategy is that it blends the strengths of data-driven and knowledge-guided design: 1) Imposing a causal structure defines interventional distributions of all variables, allowing benchmarks to ask flexible causal queries of the simulator (e.g., about mediation or path-specific effects), 2) Fitting the initial states of variables adds realistic complexity to structural equations. 3) A sequential model creates causal effects of future outcomes that are interesting functions of early interventions and context variables even when transition mechanisms are hand-designed.
@@
and aim to estimate $\mathrm{CATE}$ from observational data generated by the simulator, as well as the average treatment effect, $\mbox{ATE}\coloneqq\mathbb{E}_{Z}[\mathrm{CATE}(Z)]$.
@@ 3.1 The base data set: Adult
IncomeSCM-1.0 is based on the widely-used Adult data set (Becker and Kohavi, 1996) and its release on the UC Irvine Machine Learning Repository. [^1] https://archive.ics.uci.edu/dataset/2/adult The set contains 48 842 observations of 14 variables and a sample weight, extracted from the 1994 US Census bureau database. The original task associated with the Adult set is to predict whether an individual will earn more than \$50 000 in a year and the full variable set comprises age, workclass, education (categorical and numeric), marital-status, occupation, relationship, race, sex, capital-gains/losses (summarized as capital-net in the simulator), hours-per-week (of work), native-country and income. We ignore the sample-weight in the simulator and drop any rows containing missing values in any column, leaving 30 162 samples. Statistics for these features are presented in Table 1.
@@
As primary outcome of interest $Y$, we use the yearly income of subjects in US dollars. In the Adult data set, income is coded as a binary variable, representing whether the yearly income of a subject exceeds \$50 000, denoted here by $R\in\{0,1\}$. For IncomeSCM, we construct a continuous income variable $Y$ by first fitting a small random forest regressor $h_{R}$ to predict the original binary outcome $R$ as a function of all covariates $X$ (excluding the added variable studies). The continuous predictions $\hat{R}_{i}=h_{R}(X_{i})\in[0,1]$ of the regressor are computed for each subject $i$ and re-scaled by a constant $c$ to achieve a desired average income across the full cohort.
@@
For a causal estimation task to be realistic, the primary intervention of interest must be possible to implement in practice, for example, in an experiment (Hernán, 2019; Hernán et al., 2022). The Adult data set contains a small number of variables that could be directly intervened upon and serve as the primary intervention. Base variables age, sex, native-country, and race can be removed from consideration, as can capital-gain, capital-loss as these are outside of the control of any feasible experiment. occupation is feasible to change on a yearly basis, but most jobs are unavailable to most people. relationship and marital-status are possible to change, through e.g., divorce, but most people would not be willing to change these in order to optimize their income, let alone participate in an experiment to do so. workclass, occupation and hours-per-week are candidates that could be intervened upon a yearly basis. education is coded in Adult by the highest attained degree (e.g., Bachelor’s degree) or the number of years of study. However, in the time-scale of the simulator (one observation per year, $t=1,...,T$), an intervention of, say $A_{t}\leftarrow\mathrm{Doctorate}$ is not feasible. Instead, to showcase the blend of data-driven and knowledge-guided design, we create an intervention on studies, a constructed variable, with values “Full-time studies”, “Day course”, “Evening course” and “No studies”, representing the study activity of a person during a year. The initial state of studies is determined by a logistic regression with manually selected coefficients for age, relationship, and education, and intercepts for each of the study types. studies directly affect the education level, current income and the propensity to study in the next year, and indirectly the future income of a subject.
@@ 3.2 The causal graph
We construct a cDAG for the Adult data set (the initial-state cDAG for IncomeSCM-1.0) by assigning direct causes to each variable. age, sex, race and native-country were taken to be base variables without causal parents in the data set. education was deemed to be causal of workclass, which in turn causes occupation and marital-status. marital-status and relationship are closely linked, and the causal direction was chosen to be from the former to the latter. Finally, hours-per-week and capital-net were determined by the previous variables, and income is the furthest downstream, directly caused also by studies, the hand-designed variable. The full graph is illustrated in Figure 1 (middle-left). A very similar graph was used by Nabi and Shpitser (2018) except, in their case, marital-status was a direct cause of education, not the other way around.
@@ 3.3 The samplers
IncomeSCM implements structural equations through samplers, functions of the values of parent variables, and independent sources of noise. The variables are either numeric or categorical. For the initial states of numeric variables $V\in\mathbb{R}$, we use (with few exceptions) an additive noise model,
@@
where $f_{V}:\mathrm{\textbf{pa}}(V)\rightarrow\mathbb{R}$ predicts $V$ given its parents—$f_{V}$ is a model of the conditional mean, $\mathbb{E}[V\mid\mathrm{\textbf{pa}}(V)]$—and $\epsilon_{V}$ is a source of homoskedastic noise (e.g., Gaussian). All structural equations for initial-state variables, except for studies, have observations $(V,\mathrm{\textbf{pa}}(V))$ available (income is a special case as described above). For these, $f_{V}$ is implemented by fitting a regression (e.g., a Random forest regressor) to the observed data, and the scale of $\epsilon_{V}$ is chosen proportional to the mean squared regression error. capital-net is an exception to this form, since the majority of capital gains/losses is 0. For this, we blend a classifier which predicts non-zero values and a conditional regressor for the case where the value is non-zero.
@@
For equations with no available data, such as for studies, $f_{V}$ is designed by hand. For studies, both the initial-state distribution and transitions are hand-crafted, as described above. For income1 and the transitions $g_{V}$ of most variables $V\in\mathcal{V}$, we use a blend of hand-designed rules and fitted regressions. This allows us to generate complex dependencies between variables without observing a time-series. For example, income1 is based on a regression of the original binary variable in the Adult data set, with hand-crafted modifications depending on studies and workclass. The transition sampler of relationshipt has a fixed probability $p_{\mathrm{stay}}$, that the person will stay in their relationship, relationshipt = relationshipt-1. With probability $1-p_{\mathrm{stay}}$, relationshipt is resampled from the same fitted model as the initial state, see Figure 1 (right). In this way, only the event that the relationship is changed is hand-designed, but what it is changed to depends on data.
@@ 4 Implementation & data set
Here, parents specify the direct causes of the initial state of the variable. sampler specifies the type and parameters of the sampler. In this case, the initial state of education is distributed according to a logistic regression of age, race, sex and native-country. seq_parents_curr specifies the direct causes of the transition probability from the current time step, and seq_parents_curr causes from the previous time step. seq_sampler specifies which function determines this probability. LogisticSampler is a general sampler, that outputs the probability of the given variable as a logistic-linear function of the direct causes specified in parents. General samplers can be reused for other data sets and are defined in samplers.py.
@@ 4.1 Benchmark data set
We construct a single-time causal effect estimation benchmark around the question: {mdframed}[backgroundcolor=verylightgray,linewidth=0.5pt,leftmargin=0pt,rightmargin=0pt,innertopmargin=1em,innerbottommargin=0.5em,roundcorner=3pt, linecolor=bordercolor] What is the effect on the income $Y_{7}$ of a subject at $t=7$, six years after a year of $A_{2}=$“Full-time studies” compared to $A_{2}=$‘No studies” at $t=2$, and then proceeding as they wish?
For this task, we create a cross-sectional observational data set by first sampling from the simulator at time $t=1,...,7$ where studies are selected by a “default” observational stochastic policy as defined by the samplers in IncomeSCM. Then, time $t=1$ is used to record the previous income level income_prev and studies studies_prev, prior to intervention. All other covariates are read off at time $t_{0}=2$. The binary intervention of interest is then performed at $t_{0}=2$, with $A_{2}=1$ representing Full-time studies and $A_{2}=0$ No studies. The outcome variable $Y_{T}=$incomeT is recorded at $T=7$. We use this to sample $n=50000$ samples with random seed $s=0$. First-order covariate statistics of covariates, interventions and outcomes for this seed can be seen in Table 1.
To create ground truth for causal effect estimates, we generate two analogous sets of trajectories where the same set of subjects $i=1,...,m$ are either all assigned $A^{i}_{2}=1$ (Full-time studies) or all assigned $A^{i}_{2}=0$ (No studies), thereby recording samples of both counterfactual outcomes $Y^{i}(0),Y^{i}(1)$ of these subjects. We control that the sets of subjects are identical before both interventions at $t_{0}=2$ by setting a shared seed, $s=1$. We use a different seed from the observational data set to measure out-of-sample generalization. The resulting data set(s) of observational and two interventional samples is referred to as IncomeSCM-1.0.CATE.
@@
    CATE, conditioned only on education (not itself a valid adjustment set), represented as 16 numeric bins, using the full adjustment set from Task 1.
Task 3 is an example of when the conditioning set differs from any valid adjustment set, requiring a slightly unusual application of common causal effect estimators.
Valid adjustment sets can be derived by standard graphical criteria like the backdoor criterion (Pearl, 2009) from the edge list in Table 5. Note, as described above, that the observed intervention is performed at time $t=2$ with $t=1$ being the initial state. Thus, a minimal adjustment set for estimating the causal effect of intervening on studies2 is the set of direct causes of this variable. In IncomeSCM-1.0, studiest at times $t>0$ are directly caused by studiest-1, incomet-1, aget, sext,educationt, and relationshipt. [^2] In the implementation, the helper variable time is also passed to enable interventions at specific time steps. Expanding this set with any pre-intervention variables (such as other variables at times $t=1,2$) also results in valid adjustment sets. However, to estimate the CATE conditioned only on education2, it is not sufficient to adjust only for this variable.
Table 1: Baseline characteristics in the IncomeSCM-1.0 cohort at $t=0$ and the base data set Adult (after removing missing values), as well as in-sample fit quality. Categorical variables are reported as number (rate) and continuous variables as mean (interquartile range). The proportion of subjects with income higher than \$50 000 is higher in IncomeSCM-1.0 than in Adult since the average income was calibrated to match the US population of 2023. Base variables do not involve modeling, only computing proportions/means. The capital-net regressor is fit only to sample with non-zero result. Fit quality is reported as $R^{2}$ for continuous variables and AUC for categorical variables. The income variable is evaluated using AUC since the raw data has binary income indicators.
@@
      United-States 45519 (91.0) 27504 (91.2)
sex = Female 16206 (32.4) 9782 (32.4 $\dagger$
@@
      White 42947 (85.9) 25933 (86.0)
      Black 4677 (9.4) 2817 (9.3)
age 41.1 (32.0, 49.0) 38.4 (28.0, 47.0) $\dagger$
@@
      HS-grad 15086 (30.2) 9840 (32.6)
      Some-college 9558 (19.1) 6678 (22.1)
@@
      Prof-specialty 6968 (13.9) 4038 (13.4)
      Craft-repair 6626 (13.3) 4030 (13.4)
@@
      Husband 17263 (34.5) 12463 (41.3)
      Wife 6767 (13.5) 1406 (4.7)
capital-net 1081.1 (0.0, 0.0) 1003.6 (0.0, 0.0)   0.11*
hours-per-week 42.0 (26.0, 57.0) 40.9 (40.0, 45.0) 0.28
@@
      Full-time studies 6586 (13.2)
      No studies 25241 (50.5)
income>50K (baseline) 21167 (42.3) 7508 (24.9)
income_prev (\$1000) 60.1 (16.1, 82.9) 0.93
income (\$1000) 78.2 (34.6, 103.1)